# DeciMamba: Exploring the Length Extrapolation Potential of Mamba

### v1

@@ 4 Method: DeciMamba
In Sec. 3 we identified that for a pre-trained model, the ERF of Mamba is dictated by the context size used during training $L_{train}$. This creates a blind spot for sequences that exceed $L_{train}$, as dependencies originating in the far input segments are not captured by the model, resulting in poor length generalization. To solve this problem we propose embedding a filtering mechanism within the pre-trained Mamba layers, with no need to re-train the model. The mechanism’s core task is to reduce the amount of tokens that the S6 layer processes, and it does so by discarding tokens of fewer importance. The entire method is visualized in Fig. 7 and encapsulates three aspects: (i) Decimation Strategy, (ii) Decimation Rate, and (iii) Decimation Scope.
Figure 7: DeciMamba: (Left) Schematic overview of DeciMamba. (Middle) By carefully inspecting the recurrent view of Mamba, we revealed the implicit filtering mechanism embedded in the recurrent gate and controlled by $\Delta$. (Right) We visualize the DeciMamba model, illustrating the decimation ratio and scope. The grey lines represent the sequence length at the input and output of each layer, and layers with empirically long ERFs are decimated.
Decimation Strategy: The Role of $\Delta_{t}$  Unrolling Mamba’s recurrent rule (Eq. 12) reveals the link between the selective $\Delta_{t}$ and the token’s importance score for future tokens.
@@
When $\Delta_{t}\rightarrow 0$ the layer discards the input token and preserves the previous hidden state. When $\Delta_{t}>0$ the hidden state ’attends’ the input token (the attention is proportional to $\Delta_{t}$) and adds it to an attenuated version of the previous hidden state (this attenuation is also proportional to $\Delta_{t}$). We note that this is always the case because $\Delta_{t}\geq 0$ and $A[i,i]<0$ by design. Hence, $\Delta_{t}$ can be interpreted as the controller of the recurrent gate, determining which tokens should impact future tokens. Since Mamba ignores input tokens with small $\Delta_{t}$, it is natural to decimate them.
@@
Decimation Ratio  For each decimating layer $s$, we propose to keep the Top-$P_{s}$ tokens with the largest mean $\Delta_{t}$ values (over the channels), where $P_{s}$ decreases gradually as we go deeper into the network according to the following formula:
$P_{s}=L_{base}\cdot\beta^{s},\quad\beta\in(0,1),\quad L_{base}\in\mathbb{N}$ (13)
@@
where $\beta,L_{base}$ are hyper-parameters representing a decay factor that controls the decimation rate (as we progress in depth) and the maximal length of the sequence after the first decimating layer. Sec.5.1 further discusses the selected pooling strategy.
Decimation Scope: Layer Selection  We turn to describe the decimation scope, which defines which layers should use the embedded decimation mechanism. We begin by explaining the guiding principles behind our method, followed by the criteria that reflects those principles.
@@
Our goal is to expand the ERF of a pre-trained Mamba model. Traditionally, DL models focus on various scales of token interactions at different layers. Therefore, a natural design choice is to decimate in layers that already focus on long-range dependencies. This approach can potentially increase their ability to learn global dependencies without negatively impacting the layers that are associated with short-term features.
As a criterion for measuring the scale embedded within each layer, we measure the ERF of Mamba based on the Mamba Mean Distance defined in Eq. 10. We thus select the layers with the highest Mamba Mean Distance, with the number of selected layers being a hyper-parameter.
@@ 5 Experiments
We evaluate DeciMamba as a context-extension approach across multiple tasks. Initially, we demonstrate the long-range understanding and retrieval capabilities of our method in the NLP domain. Subsequently, we assess the retrieval abilities of our method using the Passkey test, and examine the language modeling capabilities of DeciMamba.
Document Retrieval
We demonstrate DeciMamba’s extrapolation abilities in a real-world long-range NLP task. In this task the model receives a query and $N_{docs}$ randomly assorted documents, and its objective is to return the id of the document that holds the answer to the query. Our data is sampled from SQuAD v2 Rajpurkar et al. (2018). We train both Mamba-130m and DeciMamba-130m to retrieve out of $N_{docs}=11$ documents (about 2k tokens), and evaluate their performance with $N_{docs}\in[11,300]$ (2k tokens to 60,000 tokens). The results are presented in Fig. 1. While Mamba is able to retrieve until $N_{docs}=40$ documents, it suffers from a sharp decline when the amount of documents is further extended. For $N_{docs}=50$ the input sequence length is about 10k tokens, and as can be seen in Sec. 5 (Passkey Retrieval) and in Fig. 14, Mamba-130m starts to suffer from ERFs around this region when trained on 2k tokens. On the contrary, DeciMamba is able to extrapolate to sequences that are more than x25 times longer than the ones seen during training ($\sim 250$ documents), without any additional computational resources.
@@
Passkey Retrieval. We fine-tune both Mamba-130M and DeciMamba-130M to retrieve a random 5-digit code hidden in a random location within a contiguous 2K token sample from Wiki-Text Merity et al. (2016). To test the models’ extrapolation abilities, during inference we increase the sequence lengths exponentially from 1K to 128K and record their performance for a variety of passkey locations within the context. Full implementation details can be found in Appendix A.1. As can be seen in Fig. 9, DeciMamba-130m significantly increases the extrapolation abilities of Mamba-130m from 16K to 128K, when trained on sequence lengths of 2k tokens only. Interestingly, we identify that the $\Delta_{t}$ values in Layer 16 of Mamba-130m capture the exact location of the passkey, as can be seen in Fig. 13. Using this finding, we display the distortion caused to the embedded sequence due to the ERF in Fig. 14. Tokens at locations < 10k have meaningful $\Delta_{t}$ values, yet tokens at locations > 10k have noisy $\Delta_{t}$’s, possibly due to the poor information flow caused by the ERF.
Language Modeling  Following Chen et al. (2023b); Mehta et al. (2022); Chen et al. (2023a), we evaluate our method on long-range language modeling using the PG-19 dataset in both zero-shot and fine-tuned regimes.
Zero-Shot. We test our method on the test set of PG-19 using the larger Mamba models (1.4b, 2.8b) in Fig. 9 (left). We observe that the selectivity scores learned during pre-training are quite effective and extend the context naturally without any training for both models. Furthermore, our method can maintain the low perplexity achieved in the short-context region. In the ablations section we further discuss our proposed decimation mechanism and show its benefit w.r.t other alternatives.
Fine-Tuning. We display DeciMamba’s performance on PG-19 perplexity in Fig. 6. We train both Mamba-130M and DeciMamba-130M with a sequence length of 2K and test their extrapolation abilities. The full training details can be found in Sec. A.4 in the appendix. While Mamba can extrapolate to context lengths that are at most x5 times longer than the training sequences, DeciMamba can extrapolate to sequences that are about x20 times longer, without any additional computational resources. Furthermore, we plot a lower bound by calculating the perplexity of a Mamba-130M model that was trained on the same context length it evaluates on (each point on the green curve is a different model). We observe that DeciMamba is close to the lower bound and diverges from it quite slowly, while utilizing significantly less computational resources. Nevertheless, for train sequences longer than 30K the ‘lower bound’ models reach an Out Of Memory (OOM) error (Nvidia RTX-A6000, 48GB of RAM), further emphasizing the importance of good extrapolation abilities when scaling to longer sequences.
Figure 8: Passkey Retrieval. DeciMamba’s global pooling layer adapts the input sequence’s length to the effective receptive field learned during short-sequence training. This allows the model to extrapolate to sequences that are significantly longer. Results for Mamba-130m.
Refer to caption Refer to caption
Figure 9: 0-shot perplexity over the PG-19 Test Set. (Left) Comparing Mamba with DeciMamba. (Right) An ablation studying the pooling mechanism, comparing ours (based on $\Delta_{t}$) to random or max-norm pooling.
@@
5.1 Ablations
Layer Selection  We ablate our layer selection methodology, as it has significant affect on the model’s performance (Figure 10). Each point on the curve represents the score achieved by a DeciMamba-130M model in the Passkey Retrieval task when trained on sequences of length 2K. The only difference between the models is the first decimating layer (x axis). We see that when the decimating layer is too shallow (e.g. layer 8), the model fails completely. As we increase the minimal decimating layer we observe a large increase in performance, until reaching the climax at layer 12. This result aligns with the fact that the $\Delta_{t}$ distribution is farther from 0 in Mamba’s early layers (all tokens are important). We hypothesize that the tokens still don’t have a strong representation at this stage, hence it makes less sense to decimate w.r.t their $\Delta_{t}$ value. After layer 12 the performance of DeciMamba starts to drop, and stabilizes at the vicinity of the baseline Mamba model. At this region we start decimating too late, as some long-dependency layers suffer from ERFs when processing the longer Passkey Retrieval sequences. The final layers (21-23) also have larger $\Delta_{t}$ values, so they should not be decimated as well. We hypothesize that these layers ’decode’ the processed embeddings back into token distributions, so the intermediate representations are yet again not fit for semantic decimation.
Figure 10: Layer Selection Ablation. Score achieved by DeciMamba-130M in the Passkey Retrieval task when trained on sequence lengths of 2K. All models differ by the first decimating layer (x axis) and stop decimating at layer 20. The result shows the model’s sensitivity to the selection of decimating layers.
Pooling Strategy  We show the importance of the proposed Top-$P_{s}$ pooling strategy in the following ablation study. Suppose we select a different pooling strategy, Top-$K\%$, which keeps the top $K\%$ tokens with the largest mean $\Delta_{t}$ values (over the channels dimension). We train a DeciMamba-130m model for the Passkey Retrieval task on sequences of length $L_{train}=2K$, with $K=50\%$ and decimation layers [12,…,20]. As can be seen in Fig. 11 (Appendix), Top-$K\%$ extends Mamba’s extrapolation abilities from 8K to 32K, yet does not do as well as Top-$P_{s}$ which extends to 128K. While Top-$K\%$ decimation decreases the sequence length greatly (the output is x512 times shorter), during extrapolation each SSM layer still faces input sequences that are much longer than the ones it has seen during training, as demonstrated in Fig. 12. We further explain this result; During training Layer 17 saw sequences of length 65 ($L_{train}=2K$, followed by a decimation rate of 50% in each layer from layer 12). During inference it sees sequences of lengths 65 and 1002 (where $L_{eval}=2K,32K$, left and right images, respectively). Since the sequence length at layer 17 of the latter (1K) is much longer than the one it saw during training (65), the extrapolation is limited by an ERF (right image, dashed orange shape). This shows the importance of keeping the lengths of the inputs to each decimating SSM layer similar to the ones seen during training, as done by Top-$P_{s}$.
Decimation Mechanism.  To emphasize the benefit of pooling w.r.t the values of $\Delta_{t}$, we examine two additional decimation mechanisms: max-norm decimation (tokens with maximal norm are kept) and random decimation (tokens are randomly kept). We equip Mamba-1.4b with each of the decimation mechanisms (decimation layer = 12, $L_{base}=8000$), and compare each model’s zero-shot perplexity in Fig. 9 (right). Max-norm pooling achieves a similar trend to $\Delta_{t}$ pooling (DeciMamba), yet performs somewhat worse. The result suggests that $\Delta_{t}$ is a better candidate for pooling, but also shows that the token’s norm also holds some information about its importance. We hypothesize that a method that combines both might achieve better performance than any one alone, but leave this question open for future work. Random pooling induces strong distortion to the embedded sequence, yet achieves better perplexity for longer contexts. This surprising result demonstrates how sensitive Mamba is to ERFs - in longer sequences, it is actually better to randomly decimate the sequence rather than process its full length.

### v2

@@ 4 Method: DeciMamba
In Sec. 3 we identified that for a pre-trained model, the ERF of Mamba is dictated by the context size used during training $L_{train}$. This creates a blind spot for sequences that exceed $L_{train}$, as dependencies originating in the far input segments are not captured by the model, resulting in poor length generalization. To solve this problem we propose embedding a filtering mechanism within the pre-trained Mamba layers, with no need to re-train the model. The mechanism’s core task is to reduce the amount of tokens that the S6 layer processes, and it does so by discarding tokens of fewer importance. The method is described in Fig. 2 and Alg. 1. It encapsulates three aspects: (i) Decimation Strategy, (ii) Decimation Rate, and (iii) Decimation Scope. We refer the reader to Appendix C where we further discuss each hyperparameter’s role and selection strategy.
Decimation Strategy (The Role of $\Delta_{t}$) To pool the most relevant tokens, we must assign each token a relative importance score. We select the $\Delta_{t}$ parameter as a proxy for this score and motivate our choice by explaining its operation. First, we apply the SSM parametrization to Eq. 2:
@@
Notice that when $\Delta_{t}\rightarrow 0$ the layer discards the input token and preserves the previous hidden state. When $\Delta_{t}>0$ the hidden state ‘attends’ the input token (the attendance is proportional to $\Delta_{t}$) and adds it to an attenuated version of the previous hidden state (this attenuation is also proportional to $\Delta_{t}$). We note that this is always the case because $\Delta_{t}\geq 0$ and $A[i,i]<0$ by design. Hence, $\Delta_{t}$ can be interpreted as the controller of the recurrent gate, determining which tokens should impact future tokens. The final importance score for each token is the mean value over all channels of its respective $\Delta_{t}$ parameter. Note that this selection induces a small interference in Mamba’s operation, as Mamba already attempts to ignore input tokens with small $\Delta_{t}$.
@@
Decimation Ratio  For each decimating layer $s$, we propose to keep the Top-$P_{s}$ tokens with the largest importance scores, where $P_{s}$ decreases gradually as we go deeper into the network according to the following formula:
$P_{s}=L_{base}\cdot\beta^{s},\quad\beta\in(0,1),\quad L_{base}\in\mathbb{N},$ (11)
@@
where $\beta,L_{base}$ are hyper-parameters representing a decay factor that controls the decimation rate (as we progress in depth) and the maximal length of the sequence after the first decimating layer. Sec. E.2 further discusses the selected pooling strategy.
@@
Decimation Scope (Layer Selection)  We turn to describe the decimation scope, which defines which layers should use the embedded decimation mechanism. We begin by explaining the guiding principles behind our method, followed by the criteria that reflects those principles. Our goal is to expand the ERF of a pre-trained Mamba model. Traditionally, DL models focus on various scales of token interactions at different layers. Therefore, a natural design choice is to decimate layers that already focus on long-range dependencies. This approach can potentially increase their ability to learn global dependencies without negatively impacting the layers that are associated with short-term features. As a criterion for measuring the scale embedded within each layer, we measure the ERF of Mamba based on the Mamba Mean Distance defined in Eq. 8. We thus select the layers with the highest Mamba Mean Distance, with the number of selected layers being a hyper-parameter. Note that during inference we perform decimation only in the pre-fill phase. Section E.4 demonstrates that its use in the decoding phase leads to a negligible improvement and thus not required.
Inference Efficiency Mamba can be computed in parallel with a prefix-scan or step-by-step like RNNs. For efficient inference, the context is processed in parallel mode (pre-fill), and then the response is generated via auto-regressive decoding in recurrent mode. DeciMamba performs pooling only during the pre-fill phase, leveraging the parallel mode to compute the global pooling operation efficiently. During decoding, Mamba and DeciMamba have identical forms (no pooling).
Complexity  In the parallel view of Mamba, the time complexity for processing a sequence of length L, is $O(L\log L)$. In DeciMamba, we first compute the delta-based importance score for each token, identify the top-P elements, and then apply S6 to those P elements. This results in an overall complexity of $O(L+P\log P)$. We ignore the complexity of selecting the top-P elements (which is $O(L\log P)$) since this computation does not depend on the state size N or the number of channels H, making it negligible compared to the other operations in both Mamba and DeciMamba.
Algorithm 1 Decimated SSM
1:$x:\ (B-\text{Batch siz}e,L-\text{Tokens},D-\text{Channels})$
2:$A:\ (D,N)\leftarrow Parameter$
3:$B:\ (B,L,N)\leftarrow S_{B}(x)$
4:$C:\ (B,L,N)\leftarrow S_{C}(x)$
5:$\Delta:\ (B,L,D)\leftarrow\tau\Delta(Parameter+S_{\Delta}(x))$
6:$\Delta_{P}\leftarrow\text{Top-P}(|\Delta|).\text{indices}$
7:$\Delta,A,B,C,x\leftarrow\Delta[\Delta_{P}],A[\Delta_{P}],B[\Delta_{P}],C[\Delta_{P}],x[\Delta_{P}]$
8:$A,B:\ (B,L,D,N)\leftarrow discretize(\Delta,A,B)$
9:return $y:\ (B,L,D)\leftarrow SSM(A,B,C)(x)$
@@ 5 Experiments
We evaluate DeciMamba as a context-extension approach across multiple tasks. First, we demonstrate the long-range understanding and retrieval capabilities of our method in the NLP domain. Next, we assess the retrieval capabilities of our method using the Passkey-Retrieval task and examine its language modeling capabilities over PG-19. Finally, we analyze and measure DeciMamba’s inference efficiency and provide ablation studies to support our claims.
LongBench This benchmark contains a variety of zero-shot long context tasks, such as: Single/Multi Document QA, Summarization, Coding, etc. We evaluate an instruction-tuned Mamba-2.8b (model: ‘xiuyul/mamba-2.8b-zephyr’ from huggingface) with and without DeciMamba (Zero-Shot) and report the results in Table 1. 0-4k, 4-8k, 8k+ are the scores for each context length group in LongBench-E, LB is the LongBench score, and ‘N/A’ means that the task does not exist in LongBench-E. In almost all tasks, DeciMamba improves the performance of the model for all context lengths. Most noticeable are the improvements at 4-8k, where Mamba starts suffering from ERFs, while DeciMamba extends its performance significantly. E.g. in TriviaQA we improve the LB result from 3.93 to 12.61, a performance gain of 220%, and improve each context length group by 24% (0-4k), 150% (4-8k), and 152% (8k+) respectively. Despite the improvement, when DeciMamba is applied to zero-shot scenarios it does not completely prevent degradation when the context length increases, motivating future research of this behavior. See Sec. A.3 for more details.
Table 1: LongBench. The results are for an instruction-tuned Mamba-2.8b model with and without DeciMamba (Zero-Shot). Avg Len = Average Length in words.
Type (Metric) Benchmark Avg Len Mamba DeciMamba
0-4k 4-8k 8k+ LB 0-4k 4-8k 8k+ LB
MultiDoc-QA (F1) 2wikimqa 4887 7.04 1.58 0.6 3.92 11.54 6.2 3.08 9.06
MultiDoc-QA (F1) Hotpotqa 9151 4.7 1.18 0.18 1.45 8.29 3.75 3.19 4.46
MultiDoc-QA (F1) Musique 11214 N/A N/A N/A 0.85 N/A N/A N/A 1.73
SingleDoc-QA (F1) Narrative QA 18409 N/A N/A N/A 0.87 N/A N/A N/A 1.74
SingleDoc-QA (F1) Qasper 3619 6.91 4.4 1.36 5.97 8.75 9.2 2.41 8.91
SingleDoc-QA (F1) Multifield QA 4559 18.26 5.71 2.45 11.16 26.05 13.67 4.63 18.58
Summarizing (Rouge-L) GovReport 8734 24.76 10.9 5.2 9.84 27.4 19.38 7.81 14.86
Summarizing (Rouge-L) QMSum 10614 N/A N/A N/A 8.18 N/A N/A N/A 7.08
Summarizing (Rouge-L) MultiNews 2113 24.58 9.79 5.34 23.15 25.45 17.2 4.43 24.58
Few-Shot (F1) TriviaQA 8209 10.38 5.59 4.11 3.93 12.9 14.02 10.36 12.61
Few-Shot (Rouge-L) SAMSum 6258 9.58 7.07 7.76 8.56 9.17 6.76 6.88 7.34
Few-Shot (Accuracy) TREC 5177 0.0 0.0 0.0 0.5 1.0 0.0 0.0 0.5
Code (Edit Sim) LCC 1235 8.12 5.61 4.17 8.13 9.4 14.25 7.63 8.67
Code (Edit Sim) RepoBench-p 4206 7.52 5.74 4.63 7.15 10.08 10.49 6.86 10.96
Synthetic (Accuracy) Passage Count 11141 2.0 0.0 0.0 0.0 3.0 0.0 0.0 0.5
Synthetic (Accuracy) Passage Ret. en 9289 0.0 0.0 0.0 0.0 9.0 1.0 0.0 1.5
Document Retrieval In this task the model receives a query and $N_{docs}$ randomly assorted documents, and its objective is to return the id of the document that holds the answer to the query. Our data is sampled from SQuAD v2 (rajpurkar2018squadv2). We train both Mamba-130m and DeciMamba-130m to retrieve out of $N_{docs}=11$ documents (about 2k tokens), and evaluate their performance with $N_{docs}\in[11,300]$ (2k to 60k tokens). The results are presented in Fig. 1 (left). While Mamba is able to retrieve until $N_{docs}=40$ documents, it suffers from a sharp decline when the amount of documents is further extended. On the contrary, DeciMamba is able to extrapolate to sequences that are more than x25 times longer than the ones seen during training ($\sim 250$ documents), and does so without any additional computational resources. See Sec. A.2 for more details.
@@
Passkey Retrieval. We fine-tune Mamba-130M and DeciMamba-130M to retrieve a random 5-digit code hidden in a random location within a contiguous 2K token sample from Wiki-Text (merity2016pointer). To test the models’ extrapolation abilities, during inference we increase the sequence lengths exponentially from 1K to 128K and record their performance for a variety of passkey locations within the context. Full implementation details are in Appendix A.1. As can be seen in Fig. 1 (right), DeciMamba-130m significantly increases the extrapolation abilities of Mamba-130m from 16K to 128K, when trained on sequence lengths of 2k tokens only. Interestingly, we identify that the $\Delta_{t}$ values in Layer 16 of Mamba-130m capture the exact location of the passkey, as can be seen in Fig. 11. Using this finding, we display the distortion caused to the embedded sequence due to the ERF in Fig. 12. Tokens at locations $<$ 10k have meaningful $\Delta_{t}$ values, yet tokens at locations $>$ 10k have noisy $\Delta_{t}$’s, possibly due to the poor information flow caused by the limited ERF.
Language Modeling  Following chen2023longlora; mehta2022long, we evaluate our method on long-range language modeling using the PG-19 dataset in both zero-shot and fine-tuned regimes.
Zero-Shot. We test our method on the test set of PG-19 using the larger Mamba models (1.4b, 2.8b) in Fig. 7 (right). We observe that the selectivity scores learned during pre-training are quite effective and extend the context naturally without any training for both models. Furthermore, our method can maintain the low perplexity achieved in the short-context region. In the ablations section we further discuss our proposed decimation mechanism and show its benefit w.r.t other alternatives.
Fine-Tuning. We display DeciMamba’s performance on PG-19 perplexity in Fig. 7 (left). We train both Mamba-130M and DeciMamba-130M with a sequence length of 2K and test their extrapolation abilities. The full training details can be found in Sec. A.4 in the appendix. While Mamba can extrapolate to context lengths that are at most x5 times longer than the training sequences, DeciMamba can extrapolate to sequences that are about x20 times longer, without any additional computational resources. Furthermore, we plot a lower bound by calculating the perplexity of a Mamba-130M model that was trained on the same context length it evaluates on (each point on the green curve is a different model). We observe that DeciMamba is close to the lower bound and diverges from it quite slowly, while utilizing significantly less computational resources. Nevertheless, for train sequences longer than 30K the ‘lower bound’ models reach an Out Of Memory (OOM) error (Nvidia RTX-A6000, 48GB of RAM), which is caused by Mamba’s O(L) training complexity. This further emphasizes the importance of good extrapolation abilities when scaling to longer sequences.
Figure 7: Perplexity Over PG-19 Test Set. (Left) The dashed purple line shows the train sequence length for Mamba-130m and DeciMamba-130m. Each point on the green curve shows a different Mamba-130m model trained on the respective context length ($L_{train}=L_{eval}$); for $L_{train}>30K$ an Out Of Memory (OOM) error occurs. (Right) Zero-Shot perplexity comparison.
Refer to caption Refer to caption
Figure 8: Ablation Studies. (Left) Model sensitivity to decimation layer selection. We show the score achieved by DeciMamba-130M in the Passkey Retrieval task when trained on sequences of length 2K. Each point on the graph is a model’s score, while decimating from layer ’Min Decimating Layer’ to layer 20. (Right) A Zero-Shot perplexity measurement comparing between different pooling mechanisms: ours (based on $\Delta_{t}$), random and max-norm pooling.
Comparison With Transformers. We evaluate equivalent Transformer models from the Pythia suite (biderman2023pythiasuiteanalyzinglarge) on the experiments described above. We find that vanilla Transformers have inferior length generalization abilities compared to Mamba. Full results are in Appendix B.1.
Inference Efficiency. We benchmark both DeciMamba and Mamba with a Nvidia RTX A6000 GPU and compare their inference speeds for the Passkey Retrieval task. From Table 3 it is evident that DeciMamba is twice faster than the baseline Mamba model, which is expected. This is because the decimation is performed at layer 12 (out of 24 layers in total), hence only half of the blocks process the whole context length L. The other half processes the decimated sequence (of length $P<<L$), resulting in a negligible computation time. Table 10 (Appendix) shows that per block, DeciMamba does not induce an additional computational overhead when compared to Mamba (first two rows). In addition, we can see that the computation time becomes negligible for blocks that are placed after the first DeciMamba block (rows three and four), no matter the input context length.
Table 3: Inference Efficieny. Inference time (seconds) of the entire model, benchmarked on Nvidia RTX A6000 GPU. DeciMamba is about twice as fast when compared to Mamba.
Model Context Length
8K 16K 32K 64K 128K 256K 512K
Mamba-130m 0.17 0.3 0.69 1.27 2.12 4.03 8.17
DeciMamba-130m 0.14 0.19 0.31 0.59 1.08 2.01 4.22
Ablation study The ablation studies can be found in Appendix E. Briefly, these studies examine three key aspects: layer selection, pooling strategy, and decimation mechanism. For layer selection, it seems that the choice of the first decimating layer significantly affects performance, with layer 12 yielding the best results in a Passkey Retrieval task. The pooling strategy study compares a Top-Ps approach with a Top-Kapproach, demonstrating that the former is more effective in extending the model’s extrapolation abilities. The decimation mechanism study compares various pooling methods, where $\Delta_{t}$-based pooling performs best, though max-norm pooling also shows promise. In Appendix E.4 we justify our design choice of pooling only during the pre-fill phase of inference. We demonstrate this by ablating an additional pooling mechanism during the decoding phase.